# Average-Case Deterministic Query Complexity of Boolean Functions with Fixed Weight

### v2

@@ 4.2 Lower bound
In this section, we are going to prove Theorem 1.2: $\mathrm{D}_{\mathrm{ave}}(f)\geq\log\frac{m}{\log n}-O(\log\log\frac{m}{\log n})$ for almost all functions $f\in\mathcal{B}_{n,m}$, showing Theorem 1.1 is tight up to an additive logarithmic term.
To illustrate the idea of the proof for Theorem 1.2, let us take the $\mathrm{XOR}_{n}$ function as an example. No matter which variable is queried next, the black points will be evenly partitioned, and the weight of the subfunction is exactly the half. Since the weight of the $\mathrm{XOR}_{n}$ function is $2^{n-1}$ and the algorithm cannot stop until the subfunction becomes constant, it has to query all $n$ variables on every input, which implies that $\mathrm{D}_{\mathrm{ave}}(\mathrm{XOR}_{n})=n$. Similarly, the main idea of our proof is to show almost all boolean functions have a similar property: regardless of which variable is queried next, all the black points are divided into two halves of nearly equal size. In other words, for almost all $f\in\mathcal{B}_{n,m}$, we shall prove the weight $\operatorname{wt}(f|_{P})$ is close to $\operatorname{wt}(f)\cdot 2^{-k}$ for any tree path $P$ that queries $k=\epsilon\log m$ variables.
@@
    Let $y^{(1)},\ldots,y^{(m)}\in\{0,1,\star\}^{n}$ be $m$ vectors, where, initially, all the variables are undetermined and marked by $\star$.
@@
    First, we draw $t_{0}=m$ balls, one by one, from a box with $2^{n-1}$ 00’s and $2^{n-1}$ $1$’s by sampling without replacement. Place these balls in the $i_{1}$-th position of each vector $y^{(1)},\ldots,y^{(m)}$ in order. Then, we discard the vectors in which ($(\neg v_{1})$-ball is placed. Suppose that there are $t_{1}\leq t_{0}$ vectors remaining. (Why $2^{n-2}$ instead of $2^{n-1}$? Because $\neg v_{1}$ balls have been discarded.)
@@
    Next, we draw $t_{1}$ balls from a box with $2^{n-2}$ zeros and $2^{n-2}$ ones, put them in the $i_{2}$-th position of the remaining vectors, and discard the vectors in which ($(\neg v_{2})$-ball is placed.
@@
    Repeat steps 2) and 3) for $k$ steps in total.
  5. (5)
    Output the number of the remaining vectors in the $k$-th step.
For sampling $f\in\mathcal{B}_{n,m}$ uniformly at random, a straightforward approach proceeds as follows: (1) select $m$ inputs $x^{(1)},\cdots,x^{(m)}\in\{0,1\}^{n}$ uniformly at random without replacement; (2) set $f(x^{(i)})=1$ for all $i=1,2,\ldots,m$; and (3) set the value of the other inputs to 00. This can also be done by repeatedly drawing balls without replacement and placing them into $m$ vectors $y^{(1)},\ldots,y^{(m)}\in\{0,1,\star\}^{n}$ that are initially empty until all the vectors are determined. Those vectors in which ($(\neg v_{i})$-ball is placed for some $i\in\{1,\ldots,k\}$ can be ignored, since they will not be accepted by the path $P$ and counted in the weight $\operatorname{wt}(f|_{P})$. Therefore, the above process samples $\operatorname{wt}(f|_{P})$ under the uniform distribution over $f\in\mathcal{B}_{n,m}$.
Denote by $t_{i}$ the random variable that indicates the number of the remaining vectors in the $i$-th step for $i=1,2,\ldots,k$. If the path $P$ computes $f$ (i.e., $f|_{P}\equiv c$), there should be either no vector $(t_{k}=0)$ or full of vectors ($t_{k}=2^{n-k}$) at the end; otherwise, it outputs a wrong answer. So it takes $\Omega(\log m)$ steps to discard all the $m$ vectors if $t_{i}\approx\frac{1}{2}t_{i-1}$ for all $i=1,2,\ldots,\Omega(\log m)$. Since, in each step, $t_{i}\approx\frac{1}{2}t_{i-1}$ holds with high probability by Hoeffding’s inequality, it is unlikely that a short path computes $f$. The above sketch is the proof strategy for Theorem 1.2.
@@
for every $j=1,2,\ldots,k$.
@@
Let $P$ be a decision tree path of length at most $\epsilon\log m$. For any $\delta\in(0,\frac{1}{2\epsilon\log m}]$ and $\epsilon\in(0,1)$,
$\Pr_{f\sim\mathcal{B}_{n,m}}\left[P\text{ is not }\delta\text{-parity for }f\right]<(2\epsilon\log m)\cdot\exp\left(-\frac{1}{2}\cdot\delta^{2}m^{1-\epsilon}\right).$
@@
Let $b\leq\epsilon\log m$ denote the length of $P$. The random variable $t_{j}=\operatorname{wt}(f|_{\rho_{j}})$, where $f$ is sampled from $\mathcal{B}_{n,m}$ uniformly at random. Let $X_{j,1},\ldots,X_{j,t_{j-1}}\in\{0,1\}$ be the random variables indicating whether each of the $t_{j-1}$ vectors remains in the $j$-th step, that is, whether it is in the on-set of $f|_{\rho_{j}}$. Random variables $X_{j,1},\ldots,X_{j,t_{j-1}}\in\{0,1\}$ are obtained from sampling without replacement. We have $t_{j}=X_{j,1}+\cdots+X_{j,t_{j-1}}$ and $\mathbb{E}[t_{j}]=\frac{1}{2}t_{j-1}$.
Let $\alpha=\frac{1}{2}(1-\delta)$ and $\beta=\frac{1}{2}(1+\delta)$. The probability that $P$ is not $\delta$-parity for $f$ is exactly
$\ \Pr_{f\sim\mathcal{B}_{n,m}}\left[P\text{ is not }\delta\text{-parity for }f\right]$
@@
Let $A_{j}$ denote the event $t_{j}\in\left[\alpha t_{j-1},\beta t_{j-1}\right]$ and $B_{j}$ denote the event $t_{j}\in\left[\alpha^{j}m,\beta^{j}m\right]$. By the union bound, (6) is at most
@@
If the event $A_{j}$ does not occur, we have $t_{j}>\beta t_{j-1}=\frac{1}{2}(1+\delta)t_{j-1}$ or $t_{j}<\alpha t_{j-1}=\frac{1}{2}(1-\delta)t_{j-1}$, which implies $|t_{j}-\mathbb{E}[t_{j}]|>\frac{1}{2}\delta t_{j-1}$. By Hoeffding’s inequality (Lemma 2.1),
@@
$\ \Pr_{f\sim\mathcal{B}_{n,m}}\left[P\text{ is not }\delta\text{-parity for }f\right]$
@@
A boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ is called $(t,\delta)$-parity if any decision tree path of length at most $t$ is a $\delta$-parity path (with respect to $f$).
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a $(t,\delta)$-parity function, and $\operatorname{wt}(f)\leq 2^{n-1}$. If $11\leq t\leq\log\operatorname{wt}(f)-$ 1 and $0\leq\delta\leq\frac{1}{2t}$, then $\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)\geq t$.
@@
for all $j=1,2,\ldots,k$, where $\rho_{j}$ is the restriction that fixes $x_{i_{p}}$ to $v_{p}$ for $p=1,2,\ldots,j$. Thus, we have $\frac{1}{2^{j}}(1-\delta)^{j}\operatorname{wt}(f)\leq\operatorname{wt}(f|_{\rho_{j}})\leq\frac{1}{2^{j}}(1+\delta)^{j}\operatorname{wt}(f)$. Note that (9) holds for any decision tree path of length at most $t$. So we have
@@
for any restriction $\rho$ fixing at most $t$ variables.
On one hand, by (10), we have
@@
Thus, $\operatorname{wt}(f|_{\rho})$ is less than $2^{n-j}$ and larger than $1$, which implies $f|_{\rho}$ cannot be constant for any restriction $\rho$ fixing at most $t$ variables. Therefore, we conclude the smallest certificate has size $t$ at least, i.e., $\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)\geq t$. ∎
Now, we are ready to prove Theorem 1.2.
@@
$\epsilon=1-\frac{1}{\log m}\left(\log\log n+3\log\log\frac{m}{\log n}+6\right).$
It is straightforward to prove by definition that $\mathrm{cost}(T,x)\geq\mathrm{C}_{x}(f)$ for any input $x\in\{0,1\}^{n}$ and any decision tree $T$ that computes $f$, which implies $\mathrm{D}_{\mathrm{ave}}(f)\geq\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)$. Our goal is to prove
@@
as $m\rightarrow\infty$. Since $4\log n\leq m\leq 2^{n-1}$, $m,n$ tend to infinity simultaneously,
Let $t=\epsilon\log m$ and $\delta=\frac{1}{2\log\frac{m}{\log n}}\leq\frac{1}{2\epsilon\log m}=\frac{1}{2t}$. Let $\mathrm{len}(P)$ denote the length of a tree path $P$. By Lemma 4.9, if $\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)<t$, then $f$ is not $(b,\delta)$-parity. That is, there exists a path $P$ which is not $\delta$-parity. Thus,
@@
By the union bound, (11) is at most
@@
5 Circuits
In this section, we study $\mathrm{D}_{\mathrm{ave}}(f)$ of circuits that consist of AND, OR, NOT gates with unbounded fan-in. Recall that the size of a circuit is defined as the number of the gates and that the depth is the length of the longest path between sink and source.
As a warm-up, we show an upper bound of $\mathrm{D}_{\mathrm{ave}}(F)=O(s)$ for general size-$s$ circuits $F$. The bound is tight up to a multiplicative factor, since $\mathrm{D}_{\mathrm{ave}}(\mathrm{XOR}_{n})=n$, and the function $\mathrm{XOR}_{n}$ is computable by a circuit of size $O(n)$ and depth $O(\log n)$.
Lemma 5.1.
@@ 5 Circuits
Since the average cost of evaluating each AND/OR/NOT gate does not exceed $2$ (Corollary B.2), it costs at most $2s$ on average to evaluate all $s$ gates. ∎
Let us define random restriction and criticality.
@@
A boolean function $f$ is $\lambda$-critical if and only if
@@
for every $p\in[0,1]$ and $t\in\mathbb{N}$.
The next lemma gives an upper bound on $\mathrm{D}_{\mathrm{ave}}(f)$ for any $\lambda$-critical function $f$.
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be $\lambda$-critical. Then, for any $\epsilon>0$,
$\mathrm{D}_{\mathrm{ave}}(f)\leq n\left(1-\frac{1}{(1+\epsilon)\lambda}\right)+\frac{1+\epsilon}{\epsilon^{2}}.$ (13)
@@
Let $p=\frac{1}{(1+\epsilon)\lambda}$. Since $f$ is $\lambda$-critical, we have $\Pr_{\rho\sim\mathcal{R}_{p}}[\mathrm{D}(f|_{\rho})\geq t]\leq(1+\epsilon)^{-t}$. Consider a query algorithm that queries each variable independently with probability $1-p$, and then applies a worst-case optimal query algorithm to the remaining subfunction $f|_{\rho}$. We have
$\mathrm{D}_{\mathrm{ave}}(f)$ $\leq\mathop{\mathbb{E}}_{\rho\sim\mathcal{R}_{p}}[|\operatorname{supp}(\rho)|+\mathrm{D}_{\mathrm{ave}}(f|_{\rho})]$
$\leq n(1-p)+\sum_{t=0}^{\infty}t\cdot\Pr_{\rho\sim\mathcal{R}_{p}}[\mathrm{D}(f|_{\rho})=t]$
$\leq n\left(1-p\right)+\sum_{t=0}^{\infty}\frac{t}{(1+\epsilon)^{t}}$
$=n\left(1-\frac{1}{(1+\epsilon)\lambda}\right)+\frac{1+\epsilon}{\epsilon^{2}}.$
@@
When $\epsilon=1$, one can give an alternative proof of the upper bound by applying the OS inequality $\mathrm{D}_{\mathrm{ave}}(f)\leq\log\mathrm{DT}_{\mathrm{size}}(f)$ (Theorem A.2) and the bound $\mathrm{DT}_{\mathrm{size}}(f)\leq O(2^{n(1-\frac{1}{2\lambda})})$ [Ros19], where $\mathrm{DT}_{\mathrm{size}}(f)$ denotes the decision tree size.
Remark 5.6.
Letting $\epsilon=(\frac{\lambda}{n})^{\frac{1}{3}}$, we have $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-\frac{1}{\lambda})+\Theta\left(\frac{n}{\lambda}\right)^{\frac{2}{3}}$ for $\lambda$-critical functions $f$.
Combining Lemma 5.4 with the previous results of criticality, one can derive upper bounds on $\mathrm{D}_{\mathrm{ave}}(f)$ for various circuit/formula classes. These bounds are incomparable with Lemma 5.1.
@@
For every CNF or DNF formula $F:\{0,1\}^{n}\rightarrow\{0,1\}$ of width $w$,
@@
For every CNF or DNF formula $F:\{0,1\}^{n}\rightarrow\{0,1\}$ of size $s$,
@@
For every $\mathsf{AC}^{0}$ circuit $F:\{0,1\}^{n}\rightarrow\{0,1\}$ of size $s$ and depth $d$,
@@
For every $\mathsf{AC}^{0}$ formula $F:\{0,1\}^{n}\rightarrow\{0,1\}$ of size $s$ and depth $d$,
@@
As we have discussed in the introduction, it is interesting to prove or disprove the tightness of these upper bounds. In order to provide evidence on the tightness of Corollary 5.7, 5.8, we prove Theorem 1.3 and show a DNF formula of width $w$ and size $2^{w}/w$ with $\mathrm{D}_{\mathrm{ave}}(f)=n\left(1-\frac{\log n}{\Theta(w)}\right)$.
Here we briefly explain how to construct the DNF formula. Although $\mathrm{D}_{\mathrm{ave}}(\mathrm{OR}_{n})=O(1)$ under a uniform input distribution, it costs $n(1-o(1))$ on average to determine the output of the OR function under a $p$-biased input distribution when $p=o(1/n)$ (Exercise 8.65 in [ODo14]). Our approach is to employ a biased function $g$ with $p=\Pr_{x\in\{0,1\}^{n}}[g(x)=1]$ and $\mathrm{D}_{\mathrm{ave}}(g)=n(1-o(1))$, which is given by Theorem 1.2, as a “simulator” of $p$-biased variable. Then, we show the composition $\mathrm{OR}\circ g$ is hard to query under a uniform distribution and is computable by a somewhat small DNF formula. As such, Theorem 1.3 follows.
@@
Let $s=\lceil\frac{2^{w}}{w}\rceil$ and $h=\lceil\frac{n}{w}\rceil$. Define $m=\lceil\frac{s}{h}\rceil$. Since $w\leq n$, we have
$m\leq\frac{s}{h}+1\leq\frac{\frac{2^{w}}{w}+1}{\frac{n}{w}}+1=\frac{2^{w}}{n}+\frac{w}{n}+1\leq\frac{2^{n}}{n}+2\leq 2^{n-1}$
for any $n\geq 4$. Since $w\geq\log n+\log\log n+3$, we have $m\geq 4\log n$. By Theorem 1.2, there exists $g\in\mathcal{B}_{w,m}$ such that
@@
Let $p=\frac{m}{2^{w}}$ denote the probability that the function $g$ outputs $1$. Note that $p\leq\frac{1}{n}+\frac{w}{n\cdot 2^{w}}+\frac{1}{2^{w}}\leq\frac{3}{n}$.

### v3

@@ 3.2 Lower bound
In this section, we prove Theorem 1.2, showing that Theorem 1.1 is tight up to an additive logarithmic term.
To illustrate the idea of the proof, let us take the $\mathrm{XOR}_{n}$ function as an example. Regardless of which variable is queried next, the black points are evenly partitioned, and the subfunction’s weight is exactly halved. Since $\mathrm{XOR}_{n}$ has weight $2^{n-1}$ and the algorithm must continue until the subfunction becomes constant, it must query all $n$ variables for every input. Thus, $\mathrm{D}_{\mathrm{ave}}(\mathrm{XOR}_{n})=n$. Similarly, the key idea of our proof is to show that most boolean functions exhibit a similar property: regardless of which variable is queried next, the black points are split into two roughly equal halves. In other words, for almost all $f\in\mathcal{B}_{n,m}$, where $\mathcal{B}_{n,m}=\{f:\{0,1\}^{n}\rightarrow\{0,1\}\mid\operatorname{wt}(f)=m\}$, we shall prove $\operatorname{wt}(f|_{P})$ is “close” to $2^{-k}m$ for any tree path $P$ querying $k=\epsilon\log m$ variables. (Ignoring the output of $P$, we view a tree path $P$ as a restriction, with $f|_{P}$ representing the subfunction restricted to $P$.)
Now, we explain the proof strategy in more detail. To sample $f\in\mathcal{B}_{n,m}$ uniformly, a straightforward approach proceeds as follows: (1) randomly select $m$ distinct inputs $x^{(1)},\cdots,x^{(m)}\in\{0,1\}^{n}$; (2) set $f(x^{(i)})=1$ for all $i$; and (3) set the remaining inputs to 00. This can also be done by repeatedly drawing $m$ vectors from $\{0,1\}^{n}$ without replacement and placing them into $m$ vectors $y^{(1)},\ldots,y^{(m)}\in\{0,1\}^{n}$. However, to estimate the probability that $\operatorname{wt}(f|_{P})$, where $f\in\mathcal{B}_{n,m}$, is close to $2^{-k}m$ for any length-$k$ tree path, we adopt a different sampling method.
Fix a tree path $P$, viewed as a restriction. Instead of sampling a random $f\in\mathcal{B}_{n,m}$ and then estimating $\operatorname{wt}(f|_{P})$, we choose to sample $\operatorname{wt}(f|_{P})$ directly, where $f\in\mathcal{B}_{n,m}$, using the following method.
Fix a tree path $P$ of length $k$, where
@@
and $c\in\{0,1\}$ is the output of the path. We denote the restriction $f|_{x_{1}=v_{1},\cdots,x_{k}=v_{k}}$ by $f|_{P}$. In $k$ rounds, for $j=1,\ldots,k$, we sample without replacement from a box with $2^{n-j}$ 0’s and $2^{n-j}$ 1’s; place the numbers in the $i_{j}$-th position of each vector, and discard vectors with ($(\neg v_{j})$ at $i_{j}$-th position. (We can safely discard these vectors, because they are not counted in the weight of $f|_{P}$.) At the end of $k$ rounds, $\operatorname{wt}(f|_{P})$ vectors remain.
Specifically, we sample $\operatorname{wt}(f|_{P})$ as follows, given a fixed path $P$ defined in (5), where $f\in\mathcal{B}_{n,m}$ uniformly:
@@
    Let $y^{(1)},\ldots,y^{(m)}\in\{0,1,\star\}^{n}$ be the $m$ vectors, where all elements are set to $\star$ initially.
@@
    In the first round, we sample $t_{0}=m$ numbers without replacement from a box with $2^{n-1}$ zeros and $2^{n-1}$ ones. We then assign these numbers sequentially to the positions $y_{i_{1}}^{(1)},\ldots,y_{i_{1}}^{(m)}$, that is, the $i_{1}$-th position of all the $m$ vectors. After that, discard the vectors with ($(\neg v_{1})$ at position $i_{1}$, that is, $y_{i_{1}}^{(p)}=\neg v_{1}$. Let $t_{1}$ be number of remaining vectors, where $t_{1}$ is a random variable equal to $\operatorname{wt}(f|_{x_{i_{1}}=v_{1}})$.
@@
    In the second round, we sample $t_{1}$ numbers without replacement from a box with $2^{n-2}$ zeros and $2^{n-2}$ ones, since $f|_{x_{i_{1}}=v_{1},x_{i_{2}}=0}$ and $f|_{x_{i_{1}}=v_{1},x_{i_{2}}=1}$ have $2^{n-2}$ inputs. Assign these numbers sequentially to the $i_{2}$-th position of the remaining $t_{1}$ vectors, and discard the vectors with ($(\neg v_{2})$ at position $i_{2}$, i.e., $y_{i_{2}}^{(p)}=\neg v_{2}$. Let $t_{2}$ be number of remaining vectors, where $t_{2}$ is a random variable equal to $\operatorname{wt}(f|_{x_{i_{1}}=v_{1},x_{i_{2}}=v_{2}})$.
@@
    Proceed for $k$ rounds. The number of remaining vectors $t_{k}$ is a random variable equal to $\operatorname{wt}(f|_{P})=\operatorname{wt}(f|_{x_{i_{1}}=v_{1},\ldots,x_{i_{k}}=v_{k}})$.
Recall that $t_{j}$ is the number of vectors remaining after the $j$-th round, and $t_{k}=\operatorname{wt}(f|_{P})$. If path $P$ correctly computes $f$ (on input $x\in\{0,1\}^{n}$ such that $x_{i_{1}}=v_{1}$, $\ldots$, $x_{i_{k}}=v_{k}$), then we must have $\operatorname{wt}(f|_{P})=0$ or $\operatorname{wt}(f|_{P})=2^{n-k}$. Intuitively, in each round, $t_{i}\approx\frac{1}{2}t_{i-1}$ holds with high probability by Hoeffding’s inequality, so it takes $\Omega(\log n)$ rounds to make $t_{k}=0$. Thus, it is unlikely that a “short” path computes $f$.
@@
for each $j=1,\ldots,k$.
Lemma 3.9 (Hoeffding’s inequality[Hoe63, Ser74]).
@@
Let $f:\{0,1\}^{n}\to\{0,1\}$ be a boolean function with $\operatorname{wt}(f)=m$. Let $P$ be a decision tree path of length at most $\epsilon\log m$. For any $\delta\in(0,\frac{1}{2\epsilon\log m}]$ and $\epsilon\in(0,1)$,
$\Pr_{f\sim\mathcal{B}_{n,m}}\left[P\text{ is not }\delta\text{-parity for }f\right]<2\epsilon\log m\cdot\exp\left(-\frac{1}{2}\cdot\delta^{2}m^{1-\epsilon}\right).$
@@
Let $b\leq\epsilon\log m$ denote the length of $P$. The random variable $t_{j}$ equals $\operatorname{wt}(f|_{\rho_{j}})$, where $f$ is sampled uniformly at random from $\mathcal{B}_{n,m}$. Let $X_{j,1}$, $\ldots$, $X_{j,t_{j-1}}\in\{0,1\}$ be random variables indicating whether each of the $t_{j-1}$ vectors is in the on-set of $f|_{\rho_{j}}$, i.e., whether it remains after the $j$-th round. Random variables $X_{j,1},\ldots,X_{j,t_{j-1}}\in\{0,1\}$ are obtained from $2^{n-j}$ zeros and $2^{n-j}$ ones by sampling without replacement. We have $t_{j}=X_{j,1}+\cdots+X_{j,t_{j-1}}$ and $\mathbb{E}[t_{j}]=\frac{1}{2}t_{j-1}$.
Let $\alpha=\frac{1}{2}(1-\delta)$ and $\beta=\frac{1}{2}(1+\delta)$. We have
$\ \Pr_{f\sim\mathcal{B}_{n,m}}\left[P\text{ is not }\delta\text{-parity with respect to }f\right]$
@@
Let $A_{j}$ be the event that $t_{j}\in\left[\alpha t_{j-1},\beta t_{j-1}\right]$, and let $B_{j}$ be the event $t_{j}\in\left[\alpha^{j}m,\beta^{j}m\right]$. By a union bound, (7) is at most
@@
If event $A_{j}$ does not occur, we have $t_{j}>\beta t_{j-1}=\frac{1}{2}(1+\delta)t_{j-1}$ or $t_{j}<\alpha t_{j-1}=\frac{1}{2}(1-\delta)t_{j-1}$, which implies $|t_{j}-\mathbb{E}[t_{j}]|>\frac{1}{2}\delta t_{j-1}$. By Hoeffding’s inequality (Lemma 3.9),
@@
$\ \Pr_{f\sim\mathcal{B}_{n,m}}\left[P\text{ is not }\delta\text{-parity with respect to }f\right]$
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a boolean function. The function $f$ is called $(t,\delta)$-parity if any path of length at most $t$ is a $\delta$-parity path with respect to $f$.
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a $(t,\delta)$-parity function with $\operatorname{wt}(f)\leq 2^{n-1}$ satisfying $11\leq t\leq\log\operatorname{wt}(f)-$ 1 and $\delta\leq\frac{1}{2t}$. Then, $\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)\geq t$.
@@
for $j=1,2,\ldots,k$, where $\rho_{j}$ is the restriction that fixes $x_{i_{p}}$ to $v_{p}$ for $p=1,2,\ldots,j$. Thus, we have $\frac{1}{2^{j}}(1-\delta)^{j}\operatorname{wt}(f)\leq\operatorname{wt}(f|_{\rho_{j}})\leq\frac{1}{2^{j}}(1+\delta)^{j}\operatorname{wt}(f)$. Note that (10) holds for any decision tree path of length $j$. So we have
@@
for any restriction $\rho$ fixing $j$ variables, where $j\leq t$.
On one hand, from (11), we have
@@
Thus, $\operatorname{wt}(f|_{\rho})$ is less than $2^{n-j}$ and larger than 00, which implies $f|_{\rho}$ cannot be constant for any restriction $\rho$ fixing at most $t$ variables. Therefore, we conclude $\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)\geq t$. ∎
Finally, one can prove Theorem 1.2 using Lemmas 3.10 and 3.12.
@@
$\epsilon=1-\frac{1}{\log m}\left(\log\log n+3\log\log\frac{m}{\log n}+5\right).$
Since $\mathrm{cost}(T,x)\geq\mathrm{C}_{x}(f)$ for any $x\in\{0,1\}^{n}$, $\mathrm{D}_{\mathrm{ave}}(f)\geq\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)$, Our goal is to prove
@@
as $n\rightarrow\infty$. Since $4\log n\leq m\leq 2^{n-1}$, $m,n$ tend to infinity simultaneously.
Let $t=\epsilon\log m$ and $\delta=(2\log\frac{m}{\log n})^{-1}\leq(2\epsilon\log m)^{-1}=\frac{1}{2t}$. Let $\mathrm{len}(P)$ denote the length of a tree path $P$. By Lemma 3.12, if $\min_{x\in\{0,1\}^{n}}\mathrm{C}_{x}(f)<t$, then $f$ is not $(b,\delta)$-parity. That is, there exists a path $P$ being not $\delta$-parity. Thus,
@@
By a union bound, (12) is at most
@@
4 DNFs, circuits, and formulas
In this section, we study $\mathrm{D}_{\mathrm{ave}}(f)$ of circuits that consist of AND, OR, NOT gates with unbounded fan-in.
As a warm-up, we show $\mathrm{D}_{\mathrm{ave}}(F)=O(s)$ for general size-$s$ circuits $F$. The bound is tight up to a multiplicative factor, since $\mathrm{D}_{\mathrm{ave}}(\mathrm{XOR}_{n})=n$, and $\mathrm{XOR}_{n}$ is computable by a circuit of size $O(n)$ and depth $O(\log n)$.
Proposition 4.1.
@@ 4 DNFs, circuits, and formulas
Notice that the average cost of evaluating each AND/OR/NOT gate does not exceed $2$ (Proposition 3.1). Therefore, it takes at most $2s$ queries on average to evaluate $s$ gates. ∎
@@
A boolean function $f$ is $\lambda$-critical if
@@
for any $p\in[0,1]$ and $t\in\mathbb{N}$.
The next lemma gives an upper bound on $\mathrm{D}_{\mathrm{ave}}(f)$ for $\lambda$-critical functions.
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be $\lambda$-critical. Then
$\mathrm{D}_{\mathrm{ave}}(f)\leq n\left(1-\frac{1}{\lambda}\right)+2\sqrt{\frac{n}{\lambda}}.$ (14)
@@
Let $\epsilon>0$ and $p=\frac{1}{(1+\epsilon)\lambda}$. Since $f$ is $\lambda$-critical, we have $\Pr_{\rho\sim\mathcal{R}_{p}}[\mathrm{D}(f|_{\rho})\geq t]\leq(1+\epsilon)^{-t}$. Consider a query algorithm that queries each variable independently with probability $1-p$, and then applies a worst-case optimal query algorithm to $f|_{\rho}$. We have
$\mathrm{D}_{\mathrm{ave}}(f)$ $\leq\mathop{\mathbb{E}}_{\rho\sim\mathcal{R}_{p}}[|\operatorname{supp}(\rho)|+\mathrm{D}(f|_{\rho})]$
$=n(1-p)+\sum_{t=1}^{n}\Pr_{\rho\sim\mathcal{R}_{p}}[\mathrm{D}(f|_{\rho})\geq t]$
$\leq n\left(1-p\right)+\sum_{t=0}^{\infty}\frac{1}{(1+\epsilon)^{t}}$
$=n\left(1-\frac{1}{(1+\epsilon)\lambda}\right)+\frac{1+\epsilon}{\epsilon}$
$=n\left(1-\frac{1}{\lambda}\right)+\frac{n}{\lambda}\cdot\frac{\epsilon}{1+\epsilon}+\frac{1+\epsilon}{\epsilon}.$
Let $\alpha=\frac{n}{\lambda}\geq 1$. The function $h(\epsilon)=\frac{\alpha\epsilon}{1+\epsilon}+\frac{1+\epsilon}{\epsilon}$ attains its minimum at $\epsilon=\frac{1}{\sqrt{\alpha}-1}$, where $h(\frac{1}{\sqrt{\alpha}-1})=2\sqrt{\alpha}$. Thus,
$\mathrm{D}_{\mathrm{ave}}(f)\leq n\left(1-\frac{1}{\lambda}\right)+2\sqrt{\frac{n}{\lambda}}.$
@@
Alternatively, one can prove $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-\frac{1}{2\lambda})+O(1)$ by combining the OS inequality $\mathrm{D}_{\mathrm{ave}}(f)\leq\log\mathrm{DT}_{\mathrm{size}}(f)$ [OS06] and the bound $\mathrm{DT}_{\mathrm{size}}(f)\leq O(2^{n(1-\frac{1}{2\lambda})})$ [Ros19].
By combining Lemma 4.4 with the existing bounds on criticality for CNFs, bounded-depth circuits, and formulas [Has86, Ros17, Ros19, Has14, HMS23], the following upper bounds on $\mathrm{D}_{\mathrm{ave}}(f)$ can be derived.
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be computable by a CNF/DNF of width $w$. Then
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be computable by a CNF/DNF of size $s$. Then
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be computable by a circuit of depth $d$ and size $s$. Then
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be computable by a formula of depth $d$ and size $s$. Then
@@
It is natural to ask whether the upper bounds above are tight. A positive answer would suggest that random restrictions with the same probability $p$ (as was used in the proof of the aforementioned results) are optimal. Toward this goal, we prove Theorem 1.4, which says there exists a DNF of width $w$ and size $\lceil 2^{w}/w\rceil$ such that $\mathrm{D}_{\mathrm{ave}}(f)=n(1-\frac{\log n}{\Theta(w)})$.
Here, we provide an outline of the proof and briefly explain how to find such a DNF formula. In contrast to the $O(1)$ average cost to determine the output of the OR function under a uniform input distribution (Proposition 3.1), it costs $n(1-o(1))$ on average under a $p$-biased input distribution when $p=o(1/n)$ (Exercise 8.65 in [ODo14]). Our approach is to employ a biased function $g$ (given by Theorem 1.2) with $p=\Pr_{x\in\{0,1\}^{n}}[g(x)=1]$ and $\mathrm{D}_{\mathrm{ave}}(g)=n(1-o(1))$ as a “simulator” of $p$-biased variable. Then, we show the composition $\mathrm{OR}\circ g$ is hard to query under a uniform distribution and is computable by a somewhat small DNF formula. As such, Theorem 1.4 follows.
@@
Let $m=\lceil\frac{2^{w}}{2n}\rceil$ and $h=\lfloor\frac{n}{w}\rfloor$ and $s=\lceil 2^{w}/w\rceil$. Observe that
$mh\leq\left(\frac{2^{w}}{2n}+1\right)\cdot\frac{n}{w}=\frac{2^{w}}{2w}+\frac{n}{w}\leq\left\lceil\frac{2^{w}}{w}\right\rceil=s.$
Since $n\geq w\geq 2\log n$, we have $m\leq\frac{2^{w}}{2n}+1\leq 2^{n-1}$ and $m\geq\frac{2^{w}}{2n}\geq\frac{n^{2}}{2n}=\frac{n}{2}\geq 4\log w.$ By Theorem 1.2, there exists $g\in\mathcal{B}_{w,m}$ such that
@@
Let $p=\frac{m}{2^{w}}$ denote the probability that $g$ outputs $1$. Note that $p\leq\frac{1}{2n}+\frac{1}{2^{w}}\leq\frac{1}{n}$.